40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 29 13:15:35 2009, Clara, Update, PEM, Dents = Bad?? small_dent.pngbig_dent1.pngbig_dent2.png
    Reply  Wed Jul 29 19:46:04 2009, rana, alberto, Update, PEM, Dents = Bad?? 
Message ID: 1811     Entry time: Wed Jul 29 19:46:04 2009     In reply to: 1806
Author: rana, alberto 
Type: Update 
Category: PEM 
Subject: Dents = Bad?? 
It looks like the MC2 chamber and/or stack has been jarred and shifted. Please be careful and use much less force and speed around the MC2 chamber.

My guess is that the work with the accelerometers around there had made the MC2 angle and
position change last night. The reason that we don't see it in the OSEMs so much is that its
a change in the actual stack position and tilt.

To recover, we changed the MC2 alignment bias to get the beam through the Faraday. This did NOT get
the beam back onto the right place on the MC TRANS QPD. For tonight we decided to not recenter that
since Rob might not like this position. We did, however, zero out the MC WFS and the PSL POS/ANG.

If the interferometer locking is OK tonight, then we (Steve and whoever else is here at 7 AM)
should recenter IP POS and IP ANG and also fix up the PSL POS and PSL ANG QPDs. You can see
in the attached picture that there are two problems to fix:
1) The knobs (circled in red and blue) are wrapped in foil. Why???
2) The handedness of the mirror mount with the orange arrow is wrong. This should be unmounted and clocked
by 90 deg. Right now the beam is nearly clipping on the mount. Also, we need to change the channel names
on the PSL POS (or maybe its ANG). It has the horizontal/vertical channels misnamed.
ELOG V3.1.3-